A chemist mixed 18ml of 10% sulfuric acid with water to produce a mixture of 5%. How much was mixed with the acid?
Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

A chemist mixed 18ml of 10% sulfuric acid with water to produce a mixture of 5%. How much was mixed with the acid
-----------------
The concentration of water is 0%
---
5% is the average of 10% and 0% --> equal amounts.

Answer by lwsshak3(11628)

A chemist mixed 18ml of 10% sulfuric acid with water to produce a mixture of 5%. How much was mixed with the acid?
let x=amt of water to mix with the acid
18*10%+x=(18+x)*5%
1.8+x=0.9+.05x
0.95x=0.9
x=0.947
amt of water to mix with the acid=0.947 ml